OpenStudy (anonymous):

the formula can be used to approximate the speed r, in miles per hour, of a car that has left skid marks of length l in ft how far will the car skid at60mph ?

OpenStudy (anonymous):

\[2\sqrt{5l}\]

OpenStudy (anonymous):

this is the equation for the problem

OpenStudy (anonymous):

\[2\sqrt{5l}=60\]\[\sqrt{5l}=\frac{60}{2}\]\[(\sqrt{5l})^2=30^2\]\[5l=900\]\[l=\frac{900}{5}\]\[l=180\]